OpenStudy (anonymous):

Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or a 0, in the computer's memory. A byte is 2^3 bits. A megabyte (MB) is 2^20 bytes. How many bits are in a megabyte?

OpenStudy (anonymous):

what is 2^20 * 2^3?

OpenStudy (anonymous):

simplified, that is

OpenStudy (anonymous):

2^(20+3), add exponents when you multiply a number with the same base

OpenStudy (anonymous):

ok.....I'm still not catching up here....

OpenStudy (anonymous):

\[\huge2^{20}*2^3=2^{20+3}=2^{23}\] \[\huge2^x+2^y=2^{x+y}\]

OpenStudy (anonymous):

oh ok...so the answer is 2^23?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ph ok :) thank you :)

OpenStudy (anonymous):

yw